Problem 9-3A a-b Crane Limited purchased a machine on account on April 2, 2018, at an invoice price of $325,020. On April 4,Calculate the annual depreciation and total depreciation over the assets life using: (Round the depreciation cost per unit t(2) Double-diminishing-balance method Year Opening Carrying Amount Depreciation Expense Accumulated Depreciation Carrying Amo(3) Units-of-production method Year Units-of-production Depreciation Expense Accumulated Depreciation Carrying Amount 2018 $
Answer #1

Part A

Cost of the machine = 325020+1820+8150 = $334990

Part 1

Year

Depreciable cost

Depreciation expense

Accumulated depreciation

Carrying amount

2018

253530

50706

50706

284284

2020

253530

50706

101412

233578

2021

253530

50706

152118

182872

2022

253530

50706

202824

132166

2023

253530

50706

253530

81460

Depreciable cost = cost of machine – residual value = 334990-81460 = 253530

Depreciation expense = 253530/5 = 50706

Part 2

Depreciation rate = 1/ estimated life * 2 = 1/5 * 2 = 40%

Year

Opening carrying amount

Depreciation expense

Accumulated depreciation

Carrying amount

2018

334990

133996

133996

200994

2020

200994

80398

214394

120596

2021

120596

39136

253530

81460

2022

81460

0

253530

81460

2023

81460

0

253530

81460

Part 3

Depreciation per unit = (334990-81460)/6326 = 40.08

Year

Units of production

Depreciation expense

Accumulated depreciation

Carrying amount

2018

999

40040

40040

294950

2020

1507

60401

100441

234549

2021

1436

57555

157996

176995

2022

1364

54669

212665

122325

2023

1020

40865

253530

81460

Part B

Double declining method

As in double declining method, the depreciation expense is high in the early years and due high depreciation expenses, net income decreases.